### v1

# Kalman Filter Enhanced GRPO for Reinforcement Learning-Based Language Model Reasoning

Abstract
Reward baseline is important for Reinforcement Learning (RL) algorithms to reduce variance in policy gradient estimates. Recently, for language modeling, Group Relative Policy Optimization (GRPO) is proposed to compute the advantage for each output by subtracting the mean reward, as the baseline, for all outputs in the group. However, it can lead to inaccurate advantage estimates in environments with highly noisy rewards, potentially introducing bias. In this work, we propose a model, called Kalman Filter Enhanced Group Relative Policy Optimization (KRPO), by using lightweight Kalman filtering to dynamically estimate the latent reward mean and variance. This filtering technique replaces the naive batch mean baseline, enabling more adaptive advantage normalization. Our method does not require additional learned parameters over GRPO. This approach offers a simple yet effective way to incorporate multiple outputs of GRPO into advantage estimation, improving policy optimization in settings where highly dynamic reward signals are difficult to model for language models. Through experiments and analyses, we show that using a more adaptive advantage estimation model, KRPO can improve the stability and performance of GRPO. The code is available at https://github.com/billhhh/KRPO_LLMs_RL.
1 Introduction
In reinforcement learning algorithms that directly optimize the policy via gradient estimates, including Policy Gradient (PG) (sutton1999policy, ), Actor-Critic (AC) (mnih2016asynchronous, ) and Proximal Policy Optimization (PPO) (schulman2017proximal, ), variance reduction of policy gradient estimate is critical for stable and sample-efficient training. A common approach to estimate the advantage of actions is to subtract a baseline, which can be interpreted as the expectation of returns by taking all actions for the current state, from the observed returns. By doing so, the model ideally should approximate the expected value of the return under the current policy. In many popular algorithms, such as AC and PPO, this baseline is often computed via an additionally maintained value model. But it will introduce further computational cost, especially for large model training.
To overcome this issue, Group Relative Policy Optimization (GRPO) (shao2024deepseekmath, ) is proposed to simplify the prediction of reward baselines by taking the mean reward over a batch or group of trajectories. While this approach can introduce bias if inaccurate action advantage was obtained when the reward distribution is highly non-stationary and sparse. This bias can lead to unstable updates and slower convergence, especially in settings where the reward signal evolves rapidly over time.
To address this issue, we propose a dynamic advantage estimation method based on a one-dimensional Kalman filter, named as Kalman Filter Enhanced Group Relative Policy Optimization (KRPO). Instead of relying on the group mean, we take one step ahead by treating the observed rewards as noisy measurements of a latent reward signal and use the Kalman filter to model the underlying rewards and noises. The resulting filtered baseline adapts over time, smoothing out noise while tracking underlying values and uncertainties in the environment. This enables more accurate and robust advantage estimation with minimal computational overhead. Our method integrates naturally with existing GRPO algorithm and requires no modification to the policy networks. It provides a principled way to incorporate uncertainty into advantage estimation without introducing additional learning components or tuning objectives. We evaluate our approach by comparing it to the standard group-mean approach, and show it improves training stability.
In summary, this paper makes the following contributions:
  • We introduce a lightweight Kalman filter approach for accurate advantage estimation. It can be seamlessly incorporated into the Group Relative Policy Optimization (GRPO) algorithm without introducing extra learning parameters.
  • We empirically demonstrate the proposed KRPO algorithm can receive higher rewards than its the standard group-mean counterparts across different datasets.
2 Related Work
2.1 Deep Reinforcement Learning
Reinforcement learning (RL) algorithms do not require strong supervision signals as supervised learning does, and they open another door for surpassing human performance. RL methods can be broadly divided into model-based and model-free methods. Model-based methods explicitly learn a model of the environment dynamics and use it for planning or policy improvement. These approaches can be sample-efficient but often suffer from model bias. In contrast, model-free methods directly learn a policy or value function from interaction with the environment without modeling the environment dynamics. Despite a higher sample complexity, model-free methods remain dominant in many practical applications due to their relative simplicity and robustness.
Among model-free approaches, value-based methods learn to approximate the expected return of actions under a given policy. The Deep Q-Network (DQN) (mnih2013playing, ) is a foundational algorithm that learns a Q-function using temporal difference learning following Bellman Equation. Several extensions improve its stability and performance, including Double DQN (van2016deep, ) to address Q-value overestimation; Dueling DQN (wang2016dueling, ), which decomposes Q-values into state-value and advantage terms.
Policy-based methods directly optimize the policy using gradient estimates. The Policy Gradient (PG) method (sutton1999policy, ) uses sampled trajectories to compute the gradient of the expected return. Trust Region Policy Optimization (TRPO) (schulman2015trust, ) introduces a trust region constraint to ensure stable policy updates, while the Proximal Policy Optimization Algorithm (PPO) (schulman2017proximal, ) simplifies this with a clipped surrogate objective.
Combining both policy and value learning, actor-critic methods were proposed. This family includes Advantage Actor-Critic (A2C), which uses a critic to estimate the advantage function, and its asynchronous version Asynchronous Advantage Actor-Critic (A3C) (mnih2016asynchronous, ), which improves stability and exploration with multi-thread training. PPO can also be viewed as an actor-critic method with a clipped objective. In the continuous action setting, Deep Deterministic Policy Gradient (DDPG) (lillicrap2015continuous, ) learns a deterministic actor and a critic. Twin Delayed Deep Deterministic policy gradient (TD3) (fujimoto2018addressing, ) improves DDPG by addressing function overestimation and policy update delays. Soft Actor-Critic (SAC) (haarnoja2018soft, ) introduces entropy regularization to encourage exploration and improve stability.
These Reinforcement Learning algorithms can perform well for sequential action control, e.g. playing Atari games (van2016deep, ), MoJoCo (todorov2012mujoco, ), or other different simulations (wang2021multi, ; wang2020soft, ). However, they are not specifically designed for large language models (LLMs). Reducing the variance of policy gradient estimates is important for achieving stable and sample-efficient training. One widely used method is to subtract a baseline from the observed returns. This baseline typically represents the expected return over all possible actions given the current state. With this subtraction, the resulting action advantage estimation should align more closely with the expected return under the current policy. In commonly used algorithms like Actor-Critic and Proximal Policy Optimization, this baseline is usually obtained from an auxiliary value function model. Nevertheless, maintaining and training this additional model significantly increases computation, particularly when working with large-scale models.
2.2 Reinforcement Learning to Improve Reasoning for LLMs
Learning a reinforcement learning model from human feedback has been proposed by Christiano et al. (christiano2017deep, ) back in 2017. Recent advancements in aligning large language models (LLMs) with human preferences have predominantly utilized reinforcement learning from human feedback (RLHF) (ouyang2022training, ). RLHF typically involves training a reward model based on human-labeled preferences and subsequently fine-tuning the LLM using reinforcement learning algorithms such as Proximal Policy Optimization (PPO). This approach has been instrumental in enhancing models like ChatGPT (achiam2023gpt, ) and LLaMA (touvron2023llama, ).
However, RLHF presents challenges, including the complexity of reward model training and the instability of reinforcement learning procedures. To address these issues, alternative methods have been proposed. Direct Preference Optimization (DPO) (rafailov2023direct, ) eliminates the need for an explicit reward model by directly optimizing the policy to align with human preferences using a classification loss. This simplification leads to more stable and efficient training. Kahneman-Tversky Optimization (KTO) (ethayarajh2024kto, ) incorporates principles from prospect theory to better model human decision-making biases. By aligning the training objective with human utility functions, KTO aims to improve the alignment of LLM outputs with human preferences. Reinforced Self-Training (ReST) (gulcehre2023reinforced, ) adopts an offline reinforcement learning approach, generating training data from the model’s own outputs and refining the policy without the need for online interaction or extensive human feedback. Rank Responses with Human Feedback (RRHF) (yuan2023rrhf, ) simplifies the alignment process by ranking multiple model-generated responses and optimizing the model to prefer higher-ranked outputs, reducing reliance on complex reward models. Reinforcement Learning from AI Feedback (RLAIF) (lee2023rlaif, ) replaces human feedback with evaluations from a separate LLM, enabling scalable alignment without the need for human-labeled data. Self-Play Fine-Tuning (SPIN) (chen2024self, ) leverages self-play mechanisms, allowing the model to generate and learn from its own data, progressively improving performance without additional human supervision. Group Relative Policy Optimization (GRPO) introduces a group-based approach to policy optimization, where the model’s outputs are evaluated relative to a group baseline.
These methods represent a focus towards more efficient and scalable approaches for aligning LLMs with human preferences, reducing dependence on extensive human feedback and complex reinforcement learning procedures. But there is limited exploration from the estimation of a more accurate advantage that can lift up model performance.
Therefore, in this paper, we propose Kalman Filter Enhanced Group Relative Policy Optimization (KRPO), which replaces the group-mean-based advantage estimation in GRPO with a Kalman-filter-based one. By treating observed rewards as noisy measurements of a latent reward signal, the filter adaptively estimates a baseline that tracks underlying uncertainties. This gets better advantage estimation with minimal computational cost.
3 Methodology
3.1 Problem Setting
We consider reinforcement learning where an agent interacts with an environment modeled as a Markov Decision Process (MDP). At a timestep, the agent observes a state $s$, selects an action $a$ according to a policy $\pi(a\mid s)$, receives a reward $r$. The goal is to maximize the expected cumulative return. Policy gradient methods rely on estimating the advantage function to reduce the variance of gradient updates.
In Group Relative Policy Optimization (GRPO), this framework is applied to large language models by sampling multiple responses (i.e., action trajectories) from the policy given the same input prompt. A shared baseline is computed as the average return across all sampled responses for the same prompt, and the advantage for each output is calculated by subtracting this mean. While this reduces variance, the group mean may be an inaccurate estimate of the expected reward for individual outputs, particularly when the reward distribution is noisy. This can lead to unstable or suboptimal updates during training.
3.2 Kalman-Filtered Advantage Estimation
We propose to replace the naive group mean baseline with a Kalman filter that maintains an online estimate of the reward distribution’s mean and variance. Each observed return $r_{i}$ is treated as a noisy measurement of the latent reward signal $x_{i}$.
The Kalman filter maintains a posterior estimate of the reward mean $\hat{x}_{i|i}$ and its associated uncertainty $P_{i|i}$, both updated recursively. Here, the subscripts follow standard Kalman notation: $\hat{x}_{i|i-1}$ denotes the prior estimate of $x_{i}$ given observations up to step $i-1$, while $\hat{x}_{i|i}$ is the posterior estimate after incorporating the observation $r_{i}$. Similarly, $P_{i|i-1}$ and $P_{i|i}$ denote the corresponding uncertainty (variance) estimates before and after observing $r_{i}$.
The update consists of two stages: (1) Prediction Step and (2) Update Step. For the prediction step:
$\hat{x}_{i|i-1}$ $=\hat{x}_{i-1|i-1},$ (1)
$P_{i|i-1}$ $=P_{i-1|i-1}+Q,$ (2)
For the update step:
$K_{i}$ $=\frac{P_{i|i-1}}{P_{i|i-1}+\sigma^{2}},$ (3)
$\hat{x}_{i|i}$ $=\hat{x}_{i|i-1}+K_{i}(r_{i}-\hat{x}_{i|i-1}),$ (4)
$P_{i|i}$ $=(1-K_{i})P_{i|i-1},$ (5)
where $Q$ is the process noise variance and $\sigma^{2}$ is the measurement noise variance. These are hyperparameters that control the smoothness and adaptability of the estimated baseline.
Given the posterior mean $\hat{x}_{i|i}$ and the posterior variance estimate after seeing each reward $P_{i|i}$, we define the advantage at timestep $i$ as
$A_{i}=\frac{r_{i}-\hat{x}_{i|i}}{\sqrt{P_{i|i}}+\varepsilon},$ (6)
where $\varepsilon$ is a small constant for numerical stability. This formulation has two benefits: it adaptively centers the reward using a filtered baseline, and it normalizes the advantage by the estimated uncertainty, which helps stabilize the scale of policy updates.
The filtered advantage estimate $A_{i}$ can be directly used in place of existing advantage estimates for models within the policy gradient family. In our implementation, we integrate it with GRPO by replacing the typical advantage computation with the Kalman-filtered version described above. Worth to be noted, this method does not require training a separate value network for baseline estimation, nor does it add new optimization objectives. The filter state is updated online during each policy update step, and incurs negligible additional computational cost.
3.3 KRPO Algorithm
The KRPO procedure is fully differentiable and can be implemented efficiently in the GRPO frameworks. After getting the group rewards, instead of getting the mean value from the group, we fetch the Kalman filtered baseline and the estimated uncertainty $P_{i|i}$.
By take the posterior variance estimation $P_{i|i}$ into account, the advantage can be more effective to learn. That is to say, the “good” actions with large rewards can be better enhanced and the “poor” actions with small rewards can be further punished.
Input: Reward function $\mathcal{R}$, LLM policy $\pi_{\theta}$, reference policy $\pi_{\text{ref}}$, prompt dataset $\mathcal{D}$
Hyperparameters: KL weight $\beta$, group size $n$, batch size $B$, learning rate $\eta$, clipping $\epsilon$, process noise $Q$, measurement noise $\sigma^{2}$ Initialize replay buffer $\mathcal{B}$, optimizer $\text{Adam}(\theta,\eta)$ ;
for each training step do
       for each prompt $(x,y^{*})\sim\mathcal{D}$ do
             for $i=1$ to $n$ do
                   Sample trajectory $y_{i}\sim\pi_{\theta}(\cdot|x)$ ;
                   Compute reward $r_{i}=\mathcal{R}(y_{i},y^{*})$ ;
            Use a Kalman Filter to estimate reward mean $\hat{x}_{i|i}$ and variance $P_{i|i}$ from $\{r_{i}\}_{i=1}^{n}$ ;
             Compute advantage: $A_{i}=\frac{r_{i}-\hat{x}_{i|i}}{\sqrt{P_{i|i}}+\varepsilon}$ ;
             Evaluate log-probs: $\log\pi_{\theta}(y_{i}|x),\log\pi_{\text{ref}}(y_{i}|x)$ ;
             Compute KL: $D_{\mathrm{KL}}=\log\pi_{\theta}-\log\pi_{\text{ref}}$ ;
             Store $(x,y_{i},A_{i},D_{\mathrm{KL}})$ into buffer $\mathcal{B}$ ;
      for each minibatch in $\mathcal{B}$ do
             Compute policy loss with clipping:
$\mathcal{L}=-\hat{\mathbb{E}}\left[\min\left(\frac{\pi_{\theta}}{\pi_{\theta_{\text{old}}}}A,\text{clip}\left(\frac{\pi_{\theta}}{\pi_{\theta_{\text{old}}}},1-\epsilon,1+\epsilon\right)A\right)+\alpha D_{\mathrm{KL}}\right]$
Perform gradient descent on $\mathcal{L}$ with gradient clipping ;
Algorithm 1 Kalman Filter Enhanced Group Relative Policy Optimization (KRPO)
4 Connections of Reward Mean and Kalman Filtering
The baseline used in GRPO model is the empirical mean of observed rewards. Specifically, for a group of returns $\{r_{i}\}_{i=1}^{n}$, the baseline is defined as
$\bar{r}=\frac{1}{n}\sum_{i=1}^{n}r_{i},$ (7)
This empirical mean can be interpreted as a special case of a Kalman filter with the following assumptions: (1) The process noise variance $Q\to 0$, indicating that the latent state (true reward mean) is assumed to be constant; (2) The measurement noise variance $\sigma^{2}$ is a constant; (3) no prior information for the initial state.
Specifically, consider the Kalman filter update equations as aforementioned Eq. 1 to Eq. 5. When $\sigma^{2}$ is constant and the system state does not change over time (i.e., the state transition matrix is the identity and the process noise is zero), the Kalman filter update reduces to a weighted average of the observations, where the weights are determined by the prior variance and the observation noise variance. In practice, however, when the full batch of $\{r_{i}\}$ is available, computing the sample mean corresponds to performing a batch update using all data equally. This is equivalent to a Bayesian update with all measurements weighted equally, which can be seen as a limiting form of the Kalman filter without temporal dynamics.
This observation motivates the use of a Kalman filter, which allows $Q>0$, and dynamically adjusts the influence of new observations via the gain $K_{i}$. This leads to a more adaptive and sequential-aware estimate of the reward baseline.
5 Experiments
5.1 Data
We use two datasets in our experiments, i.e. Arithmetic Dataset and OpenMathInstruct Dataset, each of which is divided into three sub-datasets: easy, normal, and hard.
Arithmetic Dataset (tinygrpo2024, ). It contains 100,000 arithmetic problems involving addition, subtraction, multiplication, and division. Each problem is represented as a natural language prompt and the target is the correct answer as a string. The arithmetic expressions include up to 6-digit numbers and up to 6 terms. We first split the 100,000 problems into a training set and a testing set using the standard 80%/20% split. During both training and testing, we filter out problems whose natural language prompt exceeds 128 characters. Then, we categorize the problems into three sub-datasets based on the number of digits and terms: the easy set includes problems with at most 3-digit numbers and at most 3 terms; the normal set includes problems with at most 5-digit numbers and at most 5 terms; and the hard set includes problems with at most 6-digit numbers and at most 6 terms.
OpenMath-Instruct Dataset (toshniwal2024openmath, ). This dataset contains 7,500 math problems from various topics, including Algebra, Geometry, Prealgebra, Precalculus, Number Theory, Intermediate Algebra, and Counting & Probability. Each problem is labeled with one of five difficulty levels provided by the dataset. We split this dataset into a training set and a testing set using the same 80%/20% ratio. As with the first dataset, we filter out problems whose question text is longer than 128 characters during training and testing. We then group the problems into three sub-datasets based on difficulty level: the easy set includes problems with level $\leq$ 2; the normal set includes problems with level $\leq$ 3; and the hard set includes problems with level $\leq$ 5.
5.2 Implementation Details
We adopt group size of 12 for both GRPO and our proposed KRPO models. Training batch size is set to 16. The model is optimized using the Adam optimizer with an initial learning rate of $5\times 10^{-6}$. To make the computation managable, we adopt Llama-3.2-1B-Instruct as our base model, to individually trained with GRPO algorithm and the proposed KRPO algorithm. By default, the Kullback–Leibler (KL) divergence term is weighted by a factor $\alpha$ of 0.01 during training; variance of the filter process noise $Q$ is set to 1e-5 and variance of the filter measurement noise $\sigma^{2}$ is set to 1e-2. For the Arithmetic dataset, the model is trained for 16,000 steps. For the OpenMathInstruct dataset, the training is conducted for 5,500 steps. For the reward design, following DeepSeek-R1-Zero model (guo2025deepseek, ) to keep simplicity, we do not use a reward model and instead assign rewards as follows: if an answer exactly equal to the ground-truth answer, the reward is 1.0; if a ground-truth answer is contained within a predicted answer, 0.5 reward is assigned; otherwise, no reward will be returned. We also encourage the model to think reasonably through “<think></think>” tags. The random seed of experiments is set to 42. All experiments are conducted on one A6000 Nvidia GPU with 48 Gigabyte of graphics memory. The proposed KRPO model and its compared models are trained and tested under the same setting to keep fair comparisons.
5.3 Experimental Results
5.3.1 Overall Performance of KRPO
Table 1: Model accuracy comparison between the KRPO with other models on Arithmetic and OpenMath-Instruct datasets. The best performance for each block with the same setting are bolded.
Dataset Models Easy Normal Hard p-value
Arithmetic Llama3.2 (touvron2023llama, ) 22.516% 7.940% 5.076% 3.6e-2
PPO (schulman2017proximal, ) 64.425% 43.684% 17.650% 2.3e-2
GRPO (shao2024deepseekmath, ) 69.468% 45.228% 18.599% 4.2e-2
KRPO (Ours) 71.764% 50.589% 20.881% -
Improvement +2.296% +5.361% +2.282% -
OpenMath-Instruct Llama3.2 (touvron2023llama, ) 13.056% 12.423% 9.717% 9.0e-3
PPO (schulman2017proximal, ) 66.871% 39.783% 27.983% 1.5e-2
GRPO (shao2024deepseekmath, ) 76.399% 54.144% 33.983% 4.7e-2
KRPO (Ours) 81.022% 66.517% 51.859% -
Improvement +4.623% +12.373% +17.876% -
We compare the accuracy performance of our proposed Kalman-filter-based Relative Policy Optimization (KRPO) against several baselines, including the original Llama3.2 model, Proximal Policy Optimization (PPO), and Group Relative Policy Optimization (GRPO). We calculated the accuracy by also assigning half correctness if the ground-truth answer is contained in the returned strings, e.g. ground-truth is “59”, but “The answer is 59” is considered half correct. Table 1 shows the results on two mathematical reasoning benchmarks: Arithmetic and OpenMath-Instruct. The results are reported across three difficulty levels: Easy, Normal, and Hard.
Across both datasets and all difficulty levels, KRPO consistently outperforms the baselines. On the Arithmetic dataset, KRPO improves over GRPO by 2.296%, 5.361%, and 2.282% on Easy, Normal, and Hard subsets, respectively. Similarly, on the OpenMath-Instruct dataset, KRPO achieves substantial gains, with improvements of 4.623%, 12.373%, and 17.876% over GRPO. These improvements suggest that using a Kalman Filter to estimate the reward baseline provides a more stable and accurate advantage estimation, which benefits policy learning, especially on more challenging problems.
Beyond the overall performance improvements, we observe interesting trends across the two datasets. We find that the raw pretrained Llama3.2 model achieves higher accuracy on Arithmetic-Easy (22.516%) compared to OpenMath-Easy (13.056%), likely due to the short and highly patterned nature of Arithmetic questions, which allows the language model to leverage surface-level cues. However, after applying reinforcement learning-based reasoning training, KRPO model trained on OpenMath outperform the accuracy on Arithmetic. This highlights that Arithmetic problems, although syntactically simpler, actually require strong numerical reasoning ability, which is the weak point of Llama3.2 models. In contrast, OpenMath problems are longer and more linguistically complex, but once the reasoning structure is learned, the model can generalize more effectively across examples.
We also show the statistical significance (p-value) of performance gaps between each model and the proposed KRPO model. All differences are statistically significant with p-values below the threshold 0.05, indicating that the observed improvements of KRPO to other models are large. These results collectively indicate that incorporating a Kalman Filter model for baseline estimation is especially beneficial in high-variance, complex reasoning environments. This aligns with the goal of reducing estimation bias and improving the stability of advantage computation during training.
5.3.2 Training Reward Curve Comparison between KRPO and GRPO
(a) Easy arithmetic questions.
(b) Normal arithmetic questions.
(c) Hard arithmetic questions.
Figure 1: The curves of returns/rewards within a batch for different difficulty levels of questions within Arithmetic Dataset.
(a) Easy math questions.
(b) Normal math questions.
(c) Hard math questions.
Figure 2: The curves of returns/rewards within a batch for different difficulty levels of questions within OpenMath-Instruct Dataset.
We also show the training return curves comparison between KRPO and GRPO on Arithmetic Dataset as Fig. 1 and on OpenMath-Instruct Dataset as Fig. 2. We use running average for reward smoothing to have a better visualization quality. We set the group size to 12 and rollout batchsize to 32. In this case, the maximal reward can be obtained at each point is 384.
On both datasets and across all difficulties, the proposed KRPO algorithm can consistently converge faster and eventually reach much higher rewards. Interestingly, on the OpenMath-Instruct dataset, the performance gap between KRPO and GRPO increases with problem difficulty. This trend suggests that KRPO provides more stable and accurate advantage estimates in harder reasoning scenarios, where variance in reward signals is likely higher and a naive baseline such as mean reward becomes less effective. This trend is also applied for the Arithmetic easy and normal subset, but not observed on the Arithmetic hard dataset. It is possibly caused by the Arithmetic hard data distribution difference with other difficulty sets.
5.4 Analysis
In this sub-section, we analyze the effect of different hyper-parameters, time efficiency and observed phenomenons. For all experiments in the analysis section are conducted on the training set of Arithmetic data.
5.4.1 KL Loss Weights $\alpha$
(a) KL divergence loss weights.
(b) Process and measure noises.
Figure 3: The training return curves of different (a) KL divergence loss weights and (b) process and measure noises.
We visualize the training curves with different KL weights in the loss $\alpha$ of KRPO by setting it to {0, 0.001, 0.01, 0.05} as Fig. 3(a). We observe that with the 0.01 $\alpha$ reaches the fastest convergence, as well as the highest returns. However, with 0.001 settings, the rewards are converged slower but eventually reaches the same point; with 0.05 and 0 settings the reward are much worse than 0.01. So $\alpha=0.01$ is set as the default setting for KRPO.
5.4.2 Model Convergence against Process Noise and Measurement Noise
We investigate the model convergence against process noise $Q$ and measurement noise $\sigma^{2}$ as Fig. 4(b). We noticed that by changing the process noise $Q$ from 1e-5 to 1e-3 and changing the measurement noise $\sigma^{2}$ from 1e-2 to 1e-1 does not change the final convergence points, just making the convergence slower. However, if make the measurement noise $\sigma^{2}$ smaller to 1e-3, the model performance will be drastically worse.
5.4.3 KL Divergence, Normalized Gradients and Reward Baseline Stabilization Analysis
(a) KL divergence between the RL trained model with the reference model.
(b) Normalized gradients comparison between GRPO and the proposed KRPO.
(c) The Kalman filtered gradually stabilize the returns to a certain level.
Figure 4: The curves of (a) KL divergence; (b) normalized gradients of GRPO and the proposed KRPO; and (c) the Kalman filter gradually stabilizes the reward baseline to a certain level.
We also examine the KL divergence of action probabilities between RL finetuned model with its reference model as Fig. 4(a). Interestingly, we notice that the KRPO has slightly higher KL divergence compared to GRPO baseline. It makes sence as we observe from Tab. 1, Llama3.2 model itself is not strong enough to perform ideally on the reasoning tasks, so deviate from the reference model but not too much would give the model more freedom to have explorations to fetch higher rewards.
For the gradients as Fig. 4, we applied gradient normalization and clipping to stabilize training by limiting the total gradient norm to a fixed $\ell_{2}$-norm threshold for the whole model. But the visualized gradient is before clipping. We can see the KRPO has higher gradient than GRPO and gradually going down as the estimated variance from Kalman Filter process can highlight the benefits / drawbacks of actions. The optimization signal is stronger and more informative, making it easier to be optimized than its baseline model.
We provide an example to show the stabilization of the baseline from the Kalman Filter in Fig. 4(c). With 12 rewards between [0,1], the Kalman filter gradually stabilizes the reward baseline.
5.4.4 Case Study
Table 2: Case study for the thinking process of GRPO and the proposed KRPO model.
Question {"type": "Algebra", "question": "If $3437^{4x}=$ 343, what is the value of $7^{4x-3}$", "expected_answer": "1"}
Model Thinking Process
GRPO First, let’s rewrite $3437^{4x}=$ 343 as $7^{4x}=7^{3}$. This is because $343=7^{3}$. By equating the exponents, we get $34x=$ 3. Solving for $x$, we find that $x=\frac{3}{4}$. Therefore, $3437^{4\cdot\frac{3}{4}}=7^{3}=$ 343 . The answer to this problem is $343$.
KRPO To solve the given equation $3437^{4x}=$ 343, we first express both sides with the same base. Since $343=7^{3}$, we can rewrite the equation as $7^{4x}=7^{3}$. This implies $34x=$ 3, so $x=\frac{3}{4}$. Now that we have the value of $x$, we can find $7^{4x-3}$ by substituting $x=\frac{3}{4}$ into the expression: $17^{4\left(\frac{3}{4}\right)-3}=7^{3-3}=7^{0}=$ 1.
We also provide a case study from OpenMath data as Tab. 2 to show the improved reasoning ability of KRPO model over GRPO baseline. From the case, we observe that GRPO can figure out the prior steps, e.g. “This implies $34x=$ 3, so $x=\frac{3}{4}$”, but it stops here and derive a wrong answer 343. On the contrary, the proposed KRPO can get the correct answer 1. For this question, the KRPO can get 12 corrects out of 12 trials, while GRPO can have 3 corrects out of 12. It is because KRPO can better award and punish the correct and incorrect answers via a more accurate advantage estimation.
6 Conclusion
In this work, we improve the model performance over GRPO by introducing KRPO, a Kalman filter-based method for accurate advantage estimation. Our approach replaces the static mean with a Kalman filtered estimate that adapts to latent reward dynamics. KRPO maintains the simplicity of GRPO and requires no modification to the policy network or additional learning parameters. Empirical results show that KRPO improves training convergence and policy performance across environments with dynamic and noisy rewards.
Appendix A Time Efficiency
We also present a comparison of the per question training time between GRPO and the proposed KRPO. The Kalman Filter is a non-parametric method. The training time is measured in seconds. According to the results, GRPO requires 2.271±0.5 seconds to complete training, while KRPO takes 2.288±0.5 seconds. The difference in training time is minimal, indicating that the introduction of the Kalman filter for baseline estimation in KRPO does not introduce much computational overhead compared to GRPO. This suggests that the improved estimation in KRPO is achieved with nearly the same training efficiency. For the testing time, the Kalman filter will not be part of the evaluation process so no time burden will be added by it.
Appendix B More Examples of Case Study
Table 3: More case study of KRPO model question answering.
Dataset Data Sample Thinking Process
Arithmetic {"question": "What is -( -985 - -604 )?", "answer": "381"} The calculation process involves: -( (-985 - (-604) ) ) = -( (-985 + 604 ) ) = -( -381 ) = 381
Arithmetic {"question": "What is 9147 + -1895 + -449?", "answer": "6803"} Calculating the expression: 9147 - 1895 = 7252, then 7252 - 449 = 6803
Arithmetic {"question": "( -74 * -25 ) + -34 = ?", "answer": "1816"} First, -74 * -25 = 1850, then 1850 + -34 = 1816
Arithmetic {"question": "Solve -( 427 - 202 ) + 514", "answer": "289"} Let’s calculate this step by step: -( 427 - 202 ) + 514 = -( 225 ) + 514, -( 225 ) = -225, -225 + 514 = 289
OpenMath {"type": "Prealgebra", "question": "Add $81.76$ to $34.587$ and round your answer to the nearest tenth.} To add 81.76 and 34.587, we first add the numbers: 81.76 + 34.587 = 116.347. Then, we round to the nearest tenth, which gives us 116.3.
OpenMath {"type": "Prealgebra", "question": "What is the greatest common factor of 32 and 48", "expected_answer": "16"} To find the greatest common factor (GCF) of 32 and 48, we need to list the factors of each number. The factors of 32 are: 1, 2, 4, 8, 16, 32. The factors of 48 are: 1, 2, 3, 4, 6, 8, 12, 16, 24, 48. The greatest common factor is the largest number that appears in both lists, which is 16.
OpenMath {"type": "Counting & Probability", "question": "Compute $\dbinom{7}{4}$.", "expected_answer": "35"} To compute $\dbinom{7}{4}$, we use the formula for combinations: $\dbinom{n}{r}=\dfrac{n!}{r!(n-r)!}$. In this case, $n=7$ and $r=4$. Plugging these values into the formula, we get: $\dbinom{7}{4}=\dfrac{7!}{4!(7-4)!}=\dfrac{7!}{4!3!}=\dfrac{7\times 6\times 5}{3\times 2\times 1}=35$.
We show more examples of how KRPO conducting the thinking process on different questions in Tab. 3. These cases show the strong reasoning ability of KRPO.
Appendix C The Analysis of Leading Statement for Language Model Reasoning
From Tab. 1 and Tab. 3 we find that compared with OpenMath-Instruct data, Arithmetic data questions is not harder but model receives much lower results, especially in the hard arithmetic questions. So we hypothesis that the leading statement might be a reason as there are leading statement in OpenMath-Instruct data, but not in Arithmetic data. So we add a leading statement "Please calculate the following expression: " to each arithmetic question.
Interestingly, by enhancing with the leading statement (denoted with KRPO*), on the hard question set, the KRPO performance has increased greatly (from 20.881% of KRPO to 40.554% of KRPO*). However, for easy and normal question sets, there are improvements with leading statements but this advantage is not as obvious as on hard question set: on easy set, KRPO 71.764% V.S. KRPO* 72.478%; on the normal set, KRPO 50.589% V.S. 51.403%. It may be caused by the learning saturation of model on easy and normal question sets as the questions have fewer terms and digits that are not that difficult to be handled; but for hard question set with a lot more terms and digits, the model will easily get lost in the understanding of just arithmetic questions.